Why MUST the Top Quark Exist?

If top doesn’t exist, the Standard Model crumples

• Masses of the W ± and Z∘ bosons cannot be explained

– Measurements of the Z boson require

Mtop = 166{+17+19 \atop −19−22} GeV/c2

• Behaviour of the bottom quark implies that it must have a partner

– Weak force predicts asymmetry in e+ e− → bb̄

• B meson “mixing” could not occur

Why are Heavy Quarks Hard to Find?

Hard to produce

• Must sift through lots of matter-antimatter collisions

• Happens about 1 in every billion collisions?

They Decay!

• A top quark decays via the weak force into a bottom quark • The bottom quark then decays into lighter quarks

Ya can’t see free quarks

• Quarks are “not allowed” to appear unescorted in public • Have to “dress” themselves up around other quarks

• Result in “jets” of particles

Dilepton Top Quark Search

t t decaying into 2 leptons create unique signature

• Two charged leptons from W ± decay

• Imbalance in total energy due to undetected neutrinos • Two jets from bottom quarks

We observe a total of 6 candidate events

• Observe 1 μ+ μ− event and 5 eμ events

• Expect 1.3 ± 0.3 events from background

Probability of background fluctuation is 3.3 × 10^{−3}

b Quark Tagging in W+Jet Events

Applied these tagging techniques to W+jet events

Sample  SVX bkg SVX tags  SLT bkg  SLT tags
W+≥ 3 jet  6.7 ± 2.1  27  15.4 ± 2.3  23

• Probability of background fluctuation is 1 × 10−6

• Agrees well with expected rate from t t production